Solve the problem. Please show steps a) Common Equity(C/E) is $400,000(No Debt), Sales are $600,000, NI is $25,000, ROE goal 15% What profit margin needed to hit goal? b) NI is $20,000. Year end assets are $300,000 Liabilities to Assets ratio is 40% What is ROE?
Answer #1

a

Profit margin% = Net income/sales
Profit margin% = 25000/600000
Profit margin% = 4.17

b

Equity to asset =1-debt to asset = 1-0.4 = 0.6

equity = asset*equity to asset = 300000*0.6=180000

ROE = Net income/total equity
ROE% = 20000/180000
ROE% = 11.11